OpenStudy (ambermarie151):

Solve the system of equations. 2x-3y=21 -6x+2y=7 Part1. Explain the steps you would take to solve the system by eliminating the x-terms . Part2. Explain the steps you would take to solve the system by eliminating the y-terms. Part3. Find the values of x and y in the system. Show you work.

OpenStudy (ambermarie151):

I'm so confused

OpenStudy (will.h):

in order to use the elimination method you have to make the variables of x or y in both equations have the same coefficient. Part1: you will need to multiply the 1st equation by 3 3(2x−3y)=21 therefore the 1st equation would turn to. 6x-9y=21 and bring the second equation −6x+2y=7 Now you may scratch both of the Xs in both equations and add the y values together and add the 21+7 together therefore we will have -7y=28 then solve for y. after that substitute y value in any of the original equations to get the x value. when doing Part 2 you'll do exactly the same of this. but for the Ys not Xs Does that help?

OpenStudy (will.h):

if you need more details let me know.

OpenStudy (will.h):

do you get this??

OpenStudy (raffle_snaffle):